OpenStudy (anonymous):

solve the system algebraically y=x^3-x^2 y=-2x^2

OpenStudy (anonymous):

Have you put y = - 2x^2 in the first equation and then tried to solve ?? If no, try it now :)

OpenStudy (anonymous):

all i can get to is -2x^2=x^3-x^2

OpenStudy (freckles):

You should add 2x^2 on both sides

OpenStudy (anonymous):

Yes and after @freckles said, take x^2 common out

OpenStudy (anonymous):

x^2+x^3?

OpenStudy (anonymous):

Yes, then take x^2 common out

OpenStudy (anonymous):

then its just x???

OpenStudy (anonymous):

It becomes x^2(1+x) = 0 Then x^2 = 0 or 1+x = 0 Now you can take it from here hopefully :)

OpenStudy (anonymous):

x= -1 y= -2?